How to graph: f(x)=2log(base 2)(x+8)-3
Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(75887)

How to graph: f(x)=2log(base 2)(x+8)-3
-----
Plot a few points and sketch a curve thru them.
If x = -6 , y = 2*log2(2)-3 = 2*1-3 = -1
---
If x = 0, y = 2*log2(8)-3 = 2*3-3 = 3
---
If x = 8, y = 2*log2(16)-3 = 2*4-3 = 5
etc.
=====================
Cheers,
Stan H.

Answer by lwsshak3(11628)

How to graph: f(x)=2log(base 2)(x+8)-3
Do it in steps:
log2(x):Basic log curve
asymptote at x=0 or y-axis
x-intercept at (1,0)
..
log(x+8): shift curve 8 units left
asymptote at x=-8
x-intercept at (-7,0)
..
log(x+8)-3: drop curve 3 units down
asymptote: no change at x=-8
x-intercept coordinates drop to (-7,3)
..
2log(base 2)(x+8)-3: slope increases slightly(final configuration)
asymptote: no change at x=-8
x-intercept coordinates remain at (-7,3)